# Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem

### v2

@@ 3.2.5 Policy
Given the current state $s_{t}$, the policy selects an action according to the probability produced by the neural network. This process continues until all customer nodes are visited and all vehicles return to the depot, concluding the episode and resulting in the construction of a valid tour. In Fig. 2, we illustrate the visiting sequence generated by the policy network for an MMHCVRP instance.
@@ 4 ECHO
The illustration of ECHO is schematically presented in Fig. 3. In this section, we first introduce the customized dual-modality node encoder and the adopted vehicle encoder for MMHCVRP, then detail the PFCA mechanism-based decoder, and finally present the proposed data augment method along with the corresponding RL training algorithm.
@@ 4.1.1 Node Encoder
To capture the local topological relationships among nodes while mitigating interference from irrelevant nodes, we propose a dual-modality node encoder, which integrates edge attributes (i.e., topological distances) with node attributes (i.e., coordinates and demands). Specifically, ECHO first exploits two FCNs to embed the initial node and edge features, respectively. Let $u_{j}=(x_{j},y_{j},d_{j}^{t})$ denote the initial attributes of the $j$th node, and let $e_{j}$ be the vector of distances from the $j$th node to all other nodes. The node and edge features are then computed as follows:
@@
$E_{j}=e_{j}\bm{W}_{ed},$ (8)
where $j\in\{0,\cdots,N\}$, and the trainable parameters include $\bm{W}_{no}\in\mathbb{R}^{3\times d}$, $\bm{W}_{ed}\in\mathbb{R}^{N\times d}$ and the deport token $\text{DT}\in\mathbb{R}^{d}$. Following the prior study [2dptr], we exploit a learnable depot token to distinguish the depot from customer nodes during training.
Then, the node encoder exploits the number of $L$ stacked Multi-Head Self-Attention Blocks to facilitate information exchange among node attributes. Let $H^{0}=\{U_{0},\cdots,U_{N}\}$, the node embedding $H^{l+1}$ is computed as follows:
@@
where BN denotes the batch normalization. The functions FF and $\text{MHA}_{\text{self}}$ denote a two-layer feed‑forward network with ReLU activation function and the multi-head self-attention mechanism, respectively. They are defined as follows:
@@
Subsequently, the node encoder exploits the vanilla cross-attention mechanism to integrate the extracted local topological features $E$ with the node embedding $H^{L}$ produced by the multi‑head self‑attention blocks.
$\mathbf{N}=H^{L}+\alpha\text{MHA}_{\text{cross}}\left(H^{L},E,E\right),$ (14)
where $\text{MHA}_{\text{cross}}$ is computed analogously to $\text{MHA}_{\text{self}}$ in Eq. (12), except that the query $Q$ is derived from $H^{L}$ while the key $K$ and value $V$ are derived from $E$. The symbol $\alpha$ denotes the gating coefficient, defined as follows:
$\alpha=\text{Sigmoid}\Bigr(\bigr(\text{MHA}_{\text{cross}}(H^{L},E,E)||H^{L}\bigr)\bm{W}_{g}\Bigr),$ (15)
where $\bm{W}_{g}\in\mathbb{R}^{2d\times 1}$. By adopting this design, ECHO efficiently integrates local topological features into the node embedding. It is worth mentioning that while certain studies [wang2024distance, reevo] also highlight the importance of local topological relationships, ECHO departs fundamentally in its methodology. Specifically, our ECHO learns to aggregate the local topological features in the embedding space, whereas these prior studies reshape attention scores among nodes during decoding using these features.
@@ 4.1.2 Vehicle Encoder
Because each vehicle’s current location corresponds to a node’s coordinates, the vehicle embedding should inherently incorporate the associated node embedding. Therefore, the vehicle encoder must fuse these node embeddings with vehicle‑specific attributes to produce the vehicle embedding. To this end, we adopt the vehicle encoder architecture proposed by the prior study [2dptr]. Specifically, similar to the node encoder, the vehicle encoder first embeds the vehicle attributes (i.e., $\chi_{i},\rho_{i},\hat{\rho}_{i}^{t},$ and $\delta^{t}_{i}$, see Table 1 for definitions about these attributes), using two fully connected layers with ReLU activations, as follows:
@@
Subsequently, the vehicle encoder applies a self‑attention layer to facilitate inter‑vehicle communication, followed by a vanilla cross‑attention layer to infuse each vehicle embedding with contextual information from node embeddings. The vehicle embedding $\mathbf{M}^{t}$ is computed as follows:
@@ 4.2 Decoder of ECHO
To efficiently select vehicle-node pairs, we design a decoder inspired by the CLIP architecture [clip]. While the prior study [2dptr] also employed a similar CLIP-based architecture, our decoder incorporates fundamentally different structural components. In particular, we propose a Parameter-Free Cross Attention (PFCA) mechanism, which explicitly amplifies the influence of the node selected at the t1t-1th time step, thereby reducing myopic decision-making.
Specifically, given the vehicle embedding $\mathbf{M}^{t}\in\mathbb{R}^{M\times d}$ and node embedding $\mathbf{N}\in\mathbb{R}^{(N+1)\times d}$ produced by their respective encoders, the proposed decoder first integrates the vehicle embedding selected at the t1t-1th time step, $\mathbf{M}^{t-1}_{s}$, with the node embedding $\mathbf{N}$ to yield the updated node embedding $\hat{\mathbf{N}}^{t}$, computed as follows:
$\hat{\mathbf{N}}^{t}=\begin{cases}\mathbf{N},&t=0,\\ \text{PFCA}(\mathbf{M}^{t-1}_{s},\mathbf{N}),&\text{otherwise},\end{cases}$ (19)
$\text{PFCA}(\mathbf{M}^{t-1}_{s},\mathbf{N})=\text{softmax}(\frac{\mathbf{N}{(\mathbf{M}^{t-1}_{s})}^{\top}}{\sqrt{d}})\mathbf{M}^{t-1}_{s}+\mathbf{N}.$ (20)
By applying the proposed PFCA mechanism, information from the previously selected vehicle is seamlessly integrated into the node embedding.
Subsequently, the vehicle embedding $\mathbf{M}^{t}$ and the updated node embedding $\hat{\mathbf{N}}^{t}$ are combined to compute the attention scores for all vehicle-node pairs at the current time step. In addition, to ensure feasibility, we mask invalid pairs by setting them to -$\infty$. For example, if a vehicle lacks sufficient remaining capacity to fulfill a node’s demand, its corresponding vehicle-node pair is masked. Formally, the attention score $\beta_{i,j}^{t}$ is defined as follows:
@@
where $\tanh(\cdot)$ denotes the hyperbolic tangent activation function, $\lambda$ denotes the hyper-parameter set identically to that in the prior study [2dptr].
Finally, the probability of selecting vehicle-node pair $(m_{i},n_{j})$ at the $t$th time step is defined as follows:
@@
After computing the probability distribution, ECHO constructs solutions for the given MMHCVRP instance using either Greedy or Sampling strategies (see Section 5.1 for more details on these strategies).
@@ 4.3 Data Augment Method
Data augment techniques have been widely applied to stabilize the RL-based training process of NCO solvers in canonical single-vehicle VRPs [pomo, sym-nco]. However, in the single‑vehicle setting, these methods need only exploit the node symmetry of the VRP. In contrast, MMHCVRP exhibits both node symmetry and vehicle‐permutation invariance. To the best of our knowledge, no prior work has simultaneously considered both intrinsic properties in solving the MMHCVRP, which may underlie their suboptimal performance.
@@
$(x,y)$ $(y,x)$
(x,1y)(x,1-y) (y,1x)(y,1-x)
(1x,y)(1-x,y) (1y,x)(1-y,x)
(1x,1y)(1-x,1-y) (1y,1x)(1-y,1-x)
As shown in Fig. 4, to address this open challenge, we propose a tailored data augment method that integrates vehicle permutation invariance and node symmetry. Specifically, following the node augment approach of prior study [pomo], we generate $K=8$ augmentations per instance by reflecting node coordinates across multiple axes, as shown in Table 2. Concurrently, for each augmented instance, we sample a random vehicle permutation of the vehicle set. Note that, since the total number of vehicle permutations $M!$ may be smaller than $K$, duplicate permutations across augmentations are unavoidable (e.g., for $M=3,3!=6<8$).
We train ECHO using the well-known REINFORCE algorithm [williams_simple_1992] with a shared baseline similar to the prior studies [pomo, sym-nco]. As detailed in Section 3.1, the objective of MMHCVRP is to minimize the maximum vehicle travel time. Accordingly, we define the total reward $R$ as the negative of this objective. The policy gradient is computed as follows:
@@
SISR $\dagger$ 6.57 - 271 10.29 - 615
GA $\dagger$ 9.21 40.18 233 15.33 48.98 479
SA $\dagger$ 7.04 7.15 130 11.13 8.16 434
AM (ICLR’19) $\dagger$ (g.) 8.49 29.22 0.08 12.68 23.23 0.14
ET (AAAI’24) $\dagger$ (g.) 7.58 15.37 0.15 11.74 14.09 0.25
DPN (ICML’24) $\dagger$ (g.) 7.50 14.16 0.18 11.54 12.15 0.30
DRL (TCYB’22) $\dagger$ (g.) 7.43 13.09 0.19 11.44 11.18 0.32
2D-Ptr (AAMAS’24) (g.) 7.20 9.59 0.04 11.12 8.07 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 7.12 8.37 0.04 10.98 6.71 0.06
ECHO (ours) (g.) 6.83 3.93 0.05 10.68 3.82 0.07
AM (ICLR’19) $\dagger$ (s.) 7.62 15.98 0.14 11.82 14.87 0.29
ET (AAAI’24) $\dagger$ (s.) 7.14 8.68 0.21 11.20 8.84 0.41
DPN (ICML’24) $\dagger$ (s.) 7.08 7.76 0.25 11.04 7.29 0.48
DRL (TCYB’22) $\dagger$ (s.) 6.97 6.09 0.30 10.90 5.93 0.60
2D-Ptr (AAMAS’24) (s.) 6.82 3.81 0.07 10.71 4.08 0.13
PARCO (NeurIPS’25) $\dagger$ (s.) 6.82 3.81 0.05 10.61 3.11 0.08
ECHO (ours) (s.) 6.64 1.01 0.28 10.44 1.50 0.49
@@
SISR $\dagger$ 4.00 - 274 6.17 - 623
GA $\dagger$ 6.89 72.25 320 10.93 77.15 623
SA $\dagger$ 4.39 9.75 289 6.80 10.21 557
AM (ICLR’19) $\dagger$ (g.) 5.51 37.75 0.08 8.10 31.28 0.13
ET (AAAI’24) $\dagger$ (g.) 4.76 19.00 0.17 7.25 17.50 0.25
DPN (ICML’24) $\dagger$ (g.) 4.60 15.00 0.19 6.94 12.48 0.40
DRL (TCYB’22) $\dagger$ (g.) 4.71 17.75 0.22 7.06 14.42 0.37
2D-Ptr (AAMAS’24) (g.) 4.48 12.00 0.05 6.75 9.40 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 4.40 10.00 0.05 6.61 7.13 0.05
ECHO (ours) (g.) 4.18 4.60 0.06 6.40 3.74 0.09
AM (ICLR’19) $\dagger$ (s.) 4.82 20.50 0.13 7.45 20.75 0.28
ET (AAAI’24) $\dagger$ (s.) 4.46 11.50 0.22 6.85 11.02 0.38
DPN (ICML’24) $\dagger$ (s.) 4.35 8.75 0.28 6.66 7.94 0.52
DRL (TCYB’22) $\dagger$ (s.) 4.34 8.50 0.36 6.65 7.78 0.76
2D-Ptr (AAMAS’24) (s.) 4.20 5.00 0.08 6.46 4.70 0.16
PARCO (NeurIPS’25) $\dagger$ (s.) 4.17 4.25 0.05 6.36 3.08 0.08
ECHO (ours) (s.) 4.03 0.96 0.37 6.24 1.12 0.51

### v3

@@ 3.2.5 Policy
Given the current state $s_{t}$, the policy selects an action according to the probability produced by the neural network. This process continues until all customer nodes are visited and all vehicles return to the depot, concluding the episode and resulting in the construction of a valid tour. In Figure 1, we illustrate the visiting sequence generated by the policy network for an MMHCVRP instance.
@@ 4 ECHO
The illustration of ECHO is schematically presented in Figure 2. In this section, we first introduce the customized dual-modality node encoder and the adopted vehicle encoder for MMHCVRP, then detail the PFCA mechanism-based decoder, and finally present the proposed data augment method along with the corresponding RL training algorithm. The characteristics of different NCO solvers are summarized in Table 2. Compared to 2D-Ptr (Liu et al., 2024) and PARCO (Berto et al., 2025), our ECHO achieves the new SOTA performance for NCO solvers (see the experimental results in Section 5). This improvement can be primarily attributed to its architectural design, which explicitly leverages edge attributes and incorporates information from historically selected vehicles, whereas existing NCO solvers tailored for MMHCVRP do not account for these aspects. The source code of ECHO is available online [^1] https://github.com/wuuu110/echo.
Table 2: Summary of NCO solvers tailored for MMHCVRP
Method Type Edge Attribute Historical Vehicle Data Augment Performance
2D-Ptr AR large gap
PARCO PAR node large gap
ECHO AR node & vehicle smallest gap
@@ 4.1.1 Node Encoder
To capture the local topological relationships among nodes while mitigating interference from irrelevant nodes, we propose a dual-modality node encoder, which integrates edge attributes (i.e., topological distances) with node attributes (i.e., coordinates and demands). Specifically, ECHO first exploits two FCNs to embed the initial node and edge features, respectively. Let $u_{j}=(x_{j},y_{j},d_{j}^{t})$ denote the initial attributes of the $j$th node. The node feature $U\in\mathbb{R}^{(N+1)\times d}$ is computed according to Eq. (7).
@@
[…]
$B^{K}=e\bm{W}^{K}_{ed},\quad B^{V}=e\bm{W}^{V}_{ed},$ (8)
$E_{i}^{K}=\sum_{j=0}^{N}A^{K}_{ij}B^{K}_{ij},\quad E_{i}^{V}=\sum_{j=0}^{N}A^{V}_{ij}B^{V}_{ij}$ (9)
$A^{K}_{ij}=\frac{\exp(C^{K}_{ij})}{\sum_{k=0}^{N}\exp(C^{K}_{ik})},\quad A^{V}_{ij}=\frac{\exp(C^{V}_{ij})}{\sum_{k=0}^{N}\exp(C^{V}_{ik})},$ (10)
$C^{K}=B^{K}\bm{W}^{K}_{s},\quad C^{V}=B^{V}\bm{W}^{V}_{s},$ (11)
where $i,j\in\{0,\cdots,N\}$, and the trainable parameters include $\bm{W}^{K}_{ed},\bm{W}^{V}_{ed}$$\bm{W}^{K}_{s}$, and $\bm{W}^{V}_{s}\in\mathbb{R}^{1\times d}$. The pseudocode of the edge feature computation process is presented in Algorithm 1. Notably, this edge embedding mechanism does not introduce parameters that scale with the number of nodes $N$. Therefore, the same set of trainable parameters can be applied to instances of different sizes without padding or a fixed maximum $N$, enabling the model to naturally generalize across problem scales. In practice, the edge features are recomputed for each instance based on its coordinates, ensuring that the encoder captures instance-specific topological relationships.
Algorithm 1 Edge Feature Computation
1:Input: Euclidean distance matrix $e$
2:Output: Edge-aware node features $E^{K}$ and $E^{V}$
[…]
4:# [batch, N+1, N+1, d] see Eq. (8)
[…]
6:# [batch, N+1, N+1] see Eq. (11)
[…]
8:# [batch, N+1, N+1] see Eq. (10)
[…]
11:# [batch, N+1, d] see Eq. (9)
12:return $E^{K},E^{V}$
Then, the node encoder exploits the number of $L$ stacked Multi-Head Self-Attention Blocks to facilitate information exchange among node attributes. Let $H^{0}=\{U_{0},\cdots,U_{N}\}$, the node embedding $H^{l+1}$ is computed according to Eq. (12).
@@
[…]
@@
[…]
$\mathbf{N}=H^{L}+\alpha\text{MHA}_{\text{cross}}\left(H^{L},E^{K},E^{V}\right),$ (17)
where $\text{MHA}_{\text{cross}}$ is computed analogously to $\text{MHA}_{\text{self}}$ in Eq. (15), except that the query $Q$ is derived from $H^{L}$ while the key $K$ and value $V$ are derived from $E^{K}$ and $E^{V}$. The symbol $\alpha$ denotes the gating coefficient, defined in Eq. (18).
$\alpha=\text{Sigmoid}\Bigr(\bigr(\text{MHA}_{\text{cross}}(H^{L},E^{K},E^{V})||H^{L}\bigr)\bm{W}_{g}\Bigr),$ (18)
[…]
@@ 4.1.2 Vehicle Encoder
Because each vehicle’s current location corresponds to a node’s coordinates, the vehicle embedding should inherently incorporate the associated node embedding. Therefore, the vehicle encoder must fuse these node embeddings with vehicle‑specific attributes to produce the vehicle embedding. To this end, we adopt the vehicle encoder architecture proposed by the prior study (Liu et al., 2024). Specifically, similar to the node encoder, the vehicle encoder first embeds the vehicle attributes (i.e., $\chi_{i},\rho_{i},\hat{\rho}_{i}^{t},$ and $\delta^{t}_{i}$, see Table 1 for definitions about these attributes), using two fully connected layers with ReLU activations, defined in Eq. (19).
@@
Subsequently, the vehicle encoder applies a self‑attention layer to facilitate inter‑vehicle communication, followed by a vanilla cross‑attention layer to infuse each vehicle embedding with contextual information from node embeddings. The vehicle embedding $\mathbf{M}^{t}$ is computed according to Eqs. (20) and (21).
@@ 4.2 Decoder of ECHO
To efficiently select vehicle-node pairs, we design a decoder inspired by the CLIP architecture (Radford et al., 2021). While the prior study (Liu et al., 2024) also employed a similar CLIP-based architecture, our decoder incorporates fundamentally different structural components. In particular, we propose a Parameter-Free Cross Attention (PFCA) mechanism, which explicitly amplifies the influence of the node selected at the t1t-1th time step, thereby reducing myopic decision-making. To the best of our knowledge, ECHO is the first work to explicitly model historically selected vehicle information to alleviate myopic decisions.
Figure 3: Illustration on the decoding processes of (a) 2D-Ptr and (b) ECHO. 2D-Ptr myopically relies on cumulative travel time to select the vehicle-node pair at each time step, which may lead to suboptimal solutions. Whereas the proposed ECHO alleviates this problem by emphasizing the vehicle selected in the preceding time step.
Specifically, given the vehicle embedding $\mathbf{M}^{t}\in\mathbb{R}^{M\times d}$ and node embedding $\mathbf{N}\in\mathbb{R}^{(N+1)\times d}$ produced by their respective encoders, the proposed decoder first integrates the vehicle embedding selected at the t1t-1th time step, $\mathbf{M}_{s}$, with the node embedding $\mathbf{N}$ to yield the updated node embedding $\hat{\mathbf{N}}^{t}$. For the initial step ($t\neq 0$), no vehicle has been selected previously; therefore, $\mathbf{M}_{s}$ is undefined and $\hat{\mathbf{N}}^{t}$ is directly set to $\mathbf{N}$. For $t>0$, $\hat{\mathbf{N}}^{t}$ is computed via the proposed PFCA mechanism, which injects the historical vehicle information into the node embeddings. Formally, $\hat{\mathbf{N}}^{t}$ is computed according to Eqs. (22) and (23).
$\hat{\mathbf{N}}^{t}=\begin{cases}\mathbf{N},&t=0,\\ \text{PFCA}(\mathbf{M}_{s},\mathbf{N}),&\text{otherwise},\end{cases}$ (22)
$\text{PFCA}(\mathbf{M}_{s},\mathbf{N})=\text{softmax}(\frac{\mathbf{N}{(\mathbf{M}_{s})}^{\top}}{\sqrt{d}})\mathbf{M}_{s}+\mathbf{N}.$ (23)
By applying the proposed PFCA mechanism, information from the previously selected vehicle is seamlessly integrated into the node embedding. The advantage of ECHO over 2D-Ptr is illustrated in Figure 3.
Subsequently, the vehicle embedding $\mathbf{M}^{t}$ and the updated node embedding $\hat{\mathbf{N}}^{t}$ are combined to compute the attention scores for all vehicle-node pairs at the current time step. In addition, to ensure feasibility, we mask invalid pairs by setting them to -$\infty$. For example, if a vehicle lacks sufficient remaining capacity to fulfill a node’s demand, its corresponding vehicle-node pair is masked. Formally, the attention score $\beta_{i,j}^{t}$ is defined according to Eq. (24).
@@
where $\tanh(\cdot)$ denotes the hyperbolic tangent activation function, $\lambda$ denotes the hyper-parameter set identically to that in the prior study (Liu et al., 2024).
Finally, the probability of selecting vehicle-node pair $(m_{i},n_{j})$ at the $t$th time step is defined in Eq. (25).
@@
After computing the probability distribution, ECHO constructs solutions for the given MMHCVRP instance using either Greedy or Sampling strategies (see Section 5.1 for more details on these strategies). Moreover, if the $i$th vehicle is selected at $t$th step, the stored previous-vehicle embedding is updated as $\mathbf{M}_{s}=\mathbf{M}^{t}_{i}$, which is then used to guide the decision in the succeeding time step. It is worth noting that the decoding strategy only affects how the vehicle–node pair is selected (i.e., argmax for greedy decoding or probabilistic sampling), whereas the update and propagation of $\mathbf{M}_{s}$ follow an identical procedure once the vehicle has been determined.
@@ 4.3 Data Augment Method
Data augment techniques have been widely applied to stabilize the RL-based training process of NCO solvers in canonical single-vehicle VRPs (Kwon et al., 2020; Kim et al., 2022). However, in the single‑vehicle setting, these methods need only exploit the node symmetry of the VRP. In contrast, MMHCVRP exhibits both node symmetry and vehicle‐permutation invariance. To the best of our knowledge, no prior work has simultaneously leveraged these two properties when solving the MMHCVRP, which may underlie their suboptimal performance.
@@
1: $(x,y)$ 2: $(y,x)$
3: (x,1y)(x,1-y) 4: (y,1x)(y,1-x)
5: (1x,y)(1-x,y) 6: (1y,x)(1-y,x)
7: (1x,1y)(1-x,1-y) 8: (1y,1x)(1-y,1-x)
As shown in Figure 4, to address this open challenge, we propose a tailored data augment method that integrates vehicle permutation invariance and node symmetry. Specifically, the augmentation is performed in two ordered steps: (1) node-level geometric reflection, followed by (2) vehicle permutation. Following the node augment approach of prior study (Kwon et al., 2020), we generate $K=8$ augmentations per instance by reflecting node coordinates across multiple axes, as shown in Table 3. During this step, only spatial coordinates are transformed, while the node demands remain unchanged and are directly replicated across the augmented instances to preserve problem feasibility.
After node reflection, we perform vehicle-level augmentation by sampling a random permutation over the vehicle indices. Each vehicle is assigned a unique identifier, and the permutation only reorders the vehicle indices while keeping their intrinsic attributes (e.g., capacities and speeds) unchanged, thereby explicitly exploiting the vehicle-permutation invariance. For example, in a case with three vehicles $m_{1},m_{2},$ and $m_{3}$, there exist $M!$ number of possible vehicle permutations, including $(m_{1},m_{2},m_{3})$, $(m_{1},m_{3},m_{2})$, $(m_{2},m_{1},m_{3})$, $(m_{2},m_{3},m_{1})$, $(m_{3},m_{2},m_{1})$, and $(m_{3},m_{1},m_{2})$. Notably, when $M!$ is smaller than $K$, duplicate vehicle permutations across augmented samples are inevitable and thus allowed. The pseudocode of the overall data augment procedure is presented in Algorithm 2.
Algorithm 2 Data Augment Procedure
1:Input: Instance $G$, augmentation factor $K=8$,
2:Output: Augmented Instance Set $\{G_{i}\}_{k=1}^{K}$
3:for $k=1$ to $K$ do
4:  Apply transformation to customers and depot coordinates according to Table 3. # The demands of Customers keep unchanged
5:  Sample a random permutation over the vehicle indices and reorder them accordingly.
6:end for
7:return $\{G_{i}\}_{k=1}^{K}$
We train ECHO using the well-known REINFORCE algorithm (Williams, 1992) with a shared baseline similar to the prior studies (Kwon et al., 2020; Kim et al., 2022). As detailed in Section 3.1, the objective of MMHCVRP is to minimize the maximum vehicle travel time. Accordingly, we define the total reward $R$ as the negative of this objective. The policy gradient is computed according to Eq. (26).
@@
5 Experimental Results
5.1 Experimental Settings
5.1.1 Datasets and Hardware configuration
Following the prior studies (Liu et al., 2024; Li et al., 2022), we assess the performance of ECHO on MMHCVRP instances with varying numbers of vehicles and nodes. Except for the cross-distribution generalization experiments in Section 5.3, the coordinates of customer nodes and depot node are uniformly sampled from the unit square $0,1]\times[0,1]$, and the demand of customer nodes are uniformly sampled from the set $\{1,2,\cdots,9\}$. For vehicles, capacities are uniformly sampled from $\{20,21,\cdots,40\}$, and speeds are uniformly sampled from the interval $[0.5,1]$. For each vehicle and node scale, the training dataset comprises the number of 1,280,000 instances, and the test dataset comprises 10,000 instances. All experiments are conducted on a computer equipped with an INTEL(R) XEON(R) GOLD 6338 CPU and two NVIDIA RTX 4090 GPUs (24GB). During inference, we employ only a single GPU.
5.1.2 Hyper-parameter
Table 4: Hyper-parameter settings
Hyperparameter Value
Training epochs 50
Batches per epoch 20,000
Batch size 64
Optimizer Adam
Initial learning rate $0.0001$
Learning rate decay 0.995
Number of attention heads 8
Embedding dimension $d$ 128
Number of encoder layers $L$ 3
Temperature parameter $\lambda$ 10
To ensure a fair comparison, we follow the hyperparameter settings of the prior study (Liu et al., 2024) for ECHO. Specifically, the model is trained for 50 epochs, with each epoch comprising 20,000 batches, and each batch containing 64 randomly generated problem instances. We employ the Adam optimizer with an initial learning rate of 0.0001 and a decay factor of 0.995 per epoch. Furthermore, gradient clipping is applied to stabilize the training process, with the maximum L2 norm set to 3.0. The number of heads in all multi-head attention is set to 8, while the embedding dimension $d$ and the number $L$ of layers in the node encoder are set to 128 and 3, respectively. The temperature parameter $\lambda$ is set to 10.
@@ 5.1.2 Hyper-parameter
SISR 6.57 - 271 10.29 - 615
GA 9.21 40.18 233 15.33 48.98 479
SA 7.04 7.15 130 11.13 8.16 434
AM (ICLR’19) (g.) 8.49 29.22 0.08 12.68 23.23 0.14
ET (AAAI’24) (g.) 7.58 15.37 0.15 11.74 14.09 0.25
DPN (ICML’24) (g.) 7.50 14.16 0.18 11.54 12.15 0.30
DRL (TCYB’22) (g.) 7.43 13.09 0.19 11.44 11.18 0.32
2D-Ptr (AAMAS’24) (g.) 7.20 9.59 0.11 11.12 8.07 0.18
PARCO (NeurIPS’25) (g.) 7.12 8.37 0.04 10.98 6.71 0.06
ECHO (ours) (g.) 6.82 3.94 0.12 10.68 3.82 0.20
AM (ICLR’19) (s.) 7.62 15.98 0.14 11.82 14.87 0.29
ET (AAAI’24) (s.) 7.14 8.68 0.21 11.20 8.84 0.41
DPN (ICML’24) (s.) 7.08 7.76 0.25 11.04 7.29 0.48
DRL (TCYB’22) (s.) 6.97 6.09 0.30 10.90 5.93 0.60
2D-Ptr (AAMAS’24) (s.) 6.82 3.81 0.13 10.71 4.08 0.22
PARCO (NeurIPS’25) (s.) 6.82 3.81 0.05 10.61 3.11 0.08
ECHO (ours) (s.) 6.63 1.00 0.40 10.44 1.48 0.96
@@
SISR 4.00 - 274 6.17 - 623
GA 6.89 72.25 320 10.93 77.15 623
SA 4.39 9.75 289 6.80 10.21 557
AM (ICLR’19) (g.) 5.51 37.75 0.08 8.10 31.28 0.13
ET (AAAI’24) (g.) 4.76 19.00 0.17 7.25 17.50 0.25
DPN (ICML’24) (g.) 4.60 15.00 0.19 6.94 12.48 0.40
DRL (TCYB’22) (g.) 4.71 17.75 0.22 7.06 14.42 0.37
2D-Ptr (AAMAS’24) (g.) 4.48 12.00 0.11 6.75 9.40 0.18
PARCO (NeurIPS’25) (g.) 4.40 10.00 0.05 6.61 7.13 0.05
ECHO (ours) (g.) 4.20 5.08 0.12 6.40 3.88 0.19
AM (ICLR’19) (s.) 4.82 20.50 0.13 7.45 20.75 0.28
ET (AAAI’24) (s.) 4.46 11.50 0.22 6.85 11.02 0.38
DPN (ICML’24) (s.) 4.35 8.75 0.28 6.66 7.94 0.52
DRL (TCYB’22) (s.) 4.34 8.50 0.36 6.65 7.78 0.76
2D-Ptr (AAMAS’24) (s.) 4.20 5.00 0.13 6.46 4.70 0.23
PARCO (NeurIPS’25) (s.) 4.17 4.25 0.05 6.36 3.08 0.08
ECHO (ours) (s.) 4.05 1.32 0.52 6.24 1.29 1.18